OpenStudy (anonymous):

Factor the algebraic expression below in terms of a single trigonometric function. cos x - sin^2 x - 1

OpenStudy (anonymous):

you can do this because of sin^2x-1 we know that cos^2x+sin^2x=1 to convert we use the sinx and 1 portion will finish on next line

OpenStudy (anonymous):

Therefore \[\sin ^2(x)-1=\cos^2(x)\] \[\cos (x)-(\cos^2(x)\]

OpenStudy (anonymous):

Does this make sense? So your single equation becomes \[\cos (x)-\cos^2(x)\]

OpenStudy (anonymous):

Do you need to take the equation further or no?

OpenStudy (anonymous):

This makes a lot more sense now, thank you. So really its all based of the Pythagorean Identity Sin^2x+Cos^2x=1?

OpenStudy (anonymous):

Exactly!